In a competitive industry consisting of 5,000 firms, the short-run marginal cost curve for each

firm is given by MC = 100 + 20Q. The demand curve faced by the industry is given as P = 500 –

0.002Q. P and MC are in $/tonne and Q is in tones.

a. Find the equilibrium price and quantity sold, for the industry and for each firm.

b. Find the producer and consumer surpluses at the equilibrium price.


1
Expert's answer
2021-12-06T17:29:11-0500

a)

MC curve for each firm is:

"MC=100+Q_s"

In a competitive market, the marginal cost curve is the supply curve.

"\therefore P=100+20Q_s"

"20Q_s=P-100"

"Q_s=\frac{P}{20}-\frac{100}{20}=\frac{P}{20}-5"

There are 5000 firms.

The industry supply curve is:

"Q_s(5000)=(\frac {P}{20}-5)\times 5000=250P-25000"

The demand curve is:

"P=500-0.002Q_D"

"\implies 0.002Q_D=500-P"

"Q_D=250000-500P"

At equilibrium:

"Q_D=Q_S"

"250000-500P=250P-25000"

"\implies P=366.66"

"Q^*=250000-500(366.66)=66670"

Therefore, the total quantity sold is 66670 tones.

Quantity sold by each firm:"=\frac{66670}{5000}=13.33" tones.


b)

Consumer surplus"=\frac{1}{2}\times( 500-366.66)\times66670"

"=4444888"

Producer surplus"=\frac{1}{2}\times(366.66-100)\times 66670"

"=8889111.1"

Producer surplus for each firm:

"=\frac{8889111.1}{5000}=1777.82"